OpenStudy (anonymous):

sqrt{-75} How would I write this in terms of i?

OpenStudy (anonymous):

Simplify the expressions below as much as possible. Leave no negative numbers under radicals and no radicals in denominators.

OpenStudy (anikay):

sqrt(15) right?

OpenStudy (anonymous):

I thought so, but the imaginary numbers thing kinda threw me off a little.

OpenStudy (anikay):

me too, them tricky problem designers

OpenStudy (anonymous):

lol ikr xD Thanks for the help.

OpenStudy (anikay):

no prob

OpenStudy (raden):

it should be -sqrt(15)

OpenStudy (anonymous):

If it were a division problem, how would i work this out? \[\sqrt{-56}\div \sqrt{-7}\]

OpenStudy (anonymous):

why would it be negative sqrt?

OpenStudy (raden):

first, u have to remember defined that sqrt(-1) = i sqrt(-3) = sqrt(-1) *sqrt(3) = i sqrt(3) sqrt(-5) = sqrt(-1) *sqrt(5) = i sqrt(5), so sqrt(-3)*sqrt(-5) = i sqrt(3) * i sqrt(5) = (i)^2 sqrt(15) = (sqrt(-1))^2 sqrt(15) = -sqrt(15)

OpenStudy (anonymous):

Thank you.

OpenStudy (raden):

do same idea for the 2nd .... welcome ^^

OpenStudy (anonymous):

please help me with this question State http://openstudy.com/study#/updates/50c7fe7de4b0a14e43688878

OpenStudy (anonymous):

since \(-75=25\times3\times (-1)\) you have \[\sqrt{-75}=\sqrt{25}\times \sqrt{3}\times \sqrt{-1}=5\sqrt{3}i\]